there are 96 black and white photographs and 160 color photographs in a scrapbook . what percent of the photographs are color ?
Answer by stanbon(75887)

there are 96 black and white photographs and 160 color photographs in a scrapbook . what percent of the photographs are color ?
-----------------
160/(160+96) = 160/256 = 62.5%
------------------
Cheers,
stan H.
--------------